OpenStudy (anonymous):

write in slope intercept point, slope -2, passes throught (-4,9)

OpenStudy (anonymous):

y=mx+b m=slope y=-2(x)+b substitute in your point 9=-2(-4)+b b=1 y=-2x+1

OpenStudy (anonymous):

line | slope 2 | 4